OpenStudy (anonymous):

what is the equation of the line through (5,1) with a slope of -3 A.y+1=3(x+5) B.y-1=3(x-5) C.y-1=-3(x-5) D.y+1=-3(x-5) ok so i got c is that the correct answer if not then what is

OpenStudy (gabebae):

Do you know how you got that answer?

OpenStudy (anonymous):

not rely im confuzzled

OpenStudy (gabebae):

The answers are in point slope form trying converting them into point slope form

OpenStudy (anonymous):

how?

OpenStudy (gabebae):

I said that wrong i'm sorry try converting them into slope intercept form

OpenStudy (gabebae):

Point slope form is y-y1= m(x-x1) Slope intercept form is y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (gabebae):

Do you know how?

OpenStudy (anonymous):

umm no

OpenStudy (gabebae):

This video might help it's been so long since I did algebra last year http://www.mathwarehouse.com/algebra/linear_equation/point-slope-to-slope-intercept-form.php

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so u know the answer

OpenStudy (gabebae):

You need to solve it first and tell me what you got so I can help

OpenStudy (gabebae):

I already got a warning earlier today I don't need another one

OpenStudy (anonymous):

yea i did i hav solved it and i got c

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

/??

OpenStudy (gabebae):

hold on

OpenStudy (triciaal):

the slope of a line will be the same for any 2 points on that line. slope is the change in y / corresponding change in x sometimes referred to as rise/run

OpenStudy (gabebae):

yup it's good

OpenStudy (anonymous):

nice so i got it right YAY

OpenStudy (triciaal):